7. Potpourri-random questions. 116 points Assume that a TCP sender A first measures the actual round trip time (RTT) as 30 msec for the irst packet, and A thus sets its estimated round trip time to 30 msec. The next actual RTT that A a. 0 msee. In response, A increases its estimated RTT to 60 msec. The next measured A compute for the next value of the estimated RTT? RTT is 80 msec. What would b. If we have packets arriving as given in the table below, provide the timing for transmitting if using Priority Scheduling and Round Robin Scheduling. Assume Red has higher priority than Green. Assume if a packet arrives at time t, the earliest it can be transmitted is time 1+1, and each packet takes one full time unit to transmit. Color Arrival Tx-priority Tx- round robin GİR 0 02 0.4 1.0 14 18 2.1 272.9 34 3.8 c. Can you experience head-of-line (HOL) blocking if using a bus for the switching fabric for a router? Why or why not?
Answer #1

a.

Formula for calculating round trip time in TCP transmission:

RoundTripEstimate = x * PreviousEstimate + (1-x) * CurrentTime

Let K be estimated Round Trip Time(RTT) previously.

After the first transmission, from the question statement, we have

30 = x * K + (1-x) * 30 ____________ (1)

After second transmission,

60 = x * 30 + (1-x) * 90 _____________(2)

Solving the (2) equation, we get

x = 0.5

For the third transmission, we have the following equation

EstimateRTT = x * 60 + (1-x) * 80

Using value of x = 0.5, we get

EstimateRTT = 70 ms

b.

Transmission times are as follows:

Color G R G G R G G G R R G
Arrival 0 0.2 0.4 1.0 1.4 1.8 2.1 2.7 2.9 3.4 3.8
Tx - Priority 1.0 2.0 6.0 7.0 3.0 8.0 9.0 10.0 4.0 5.0 11.0
Tx - Round Robin 1.0 2.0 3.0 4.0 5.0 6.0 7.0 8.0 9.0 10.0 11.0

c.

When some packet which is being transmitted, blocks network traffic going to some other destination from being transmitted, there is head-of-line blocking. If bus is used for the switching fabric in a router, we can experience head-of-line blocking. Since only one packet can go through the bus at one time, other packets waiting to transmit will have to wait till the current packet is fully transmitted. Thus, head-of-line blocking is experienced.